# An Infinite Transitivity Theorem

### v3

@@ 1.2. Outline
Let $A$ be a nonunital, $\sigma$-unital $C^{*}$-algebra, and $\pi:\mathscr{Q}(A)\rightarrow\mathscr{B}(H)$ a direct sum of countably many inequivalent P-representations $\pi_{j}:\mathscr{Q}(A)\rightarrow\mathscr{B}(H_{j})$ and $K_{j}$ a separable subspace of $H_{j}$. Then the map $\mathscr{Q}(A)\rightarrow\bigoplus_{j}\mathscr{B}(K_{j})$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $K=\bigoplus_{j}K_{j}$.
As in theorem 3, we can choose $c$ as we like. With the theory complete, under the existence of a P-point, we produce a P-state, on the corona algebra $\mathscr{Q}(A)$ of a nonunital $A$ with a $\sigma$-unit of projections. However, as implicitly shown in [BI13], there is a forcing in which these states cannot exist.
@@ 2.2. Excision, P-Excision, and Quantum P-points
A P-filter is a poset $(\mathscr{F},\leq)$ which is upwards closed and $\sigma$-downwards directed. A P-point is an ultrafilter which is a P-filter under $\subseteq^{*}$, inclusion modulo finite sets. A quantum P-point is a maximal quantum filter which is P-filter under $\ll$. That is, for any countable family $\mathscr{F}\subseteq\mathscr{F}_{\varphi}$, there is $a\in\mathscr{F}_{\varphi}$ such that $a\ll b$ for all $b\in\mathscr{F}$. We call $\varphi$ a P-state if $\mathscr{F}_{\varphi}$ is a quantum P-point. Likewise, we call a representation $\pi$ of $C$ a P-representation if it is equivalent to the GNS-representation corresponding to a P-state.
@@ 2.3. Equivalences
A partially ordered set $(\mathscr{F},\leq)$ is countably closed if every decreasing sequence in $\mathscr{F}$ has a lower bound. A subset $\mathscr{P}$ of $\mathscr{F}$ is cofinal if for every $a\in\mathscr{F}$, there exists $b\in\mathscr{P}$ such that $b\leq a$. Lastly, a totally ordered set $(\mathscr{P},<)$ with no endpoints is an $\eta_{1}$-set if for all countable $X,Y\subseteq\mathscr{P}$ such that $a<b$ for all $a\in X$, $b\in Y$, there exists $c\in\mathscr{P}$ s.t. $a<c<b$ for all $a\in X$, $b\in Y$.
@@
If $C$ has real rank zero, $\varphi$ is a state on $C$, and $(C,\varphi)$ is countably degree-$1$ saturated, then $\mathscr{P}_{\varphi}$ is downwards cofinal in $\mathscr{F}_{\varphi}$.
@@
By [FA19, 5.3.10], for each $\epsilon\in(0,1)$, there exists $q\in\mathscr{P}_{\varphi}$ s.t. $q\leq a+\epsilon$, so
$\|aq-q\|^{2}\leq\|(qa-q)^{*}(aq-q)\|=\|q(a-1)^{2}q\|\leq\|qaq-q\|\leq\epsilon.$
Therefore $t(x)$ is satisfiable, so realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$. Then we repeat this with $t^{\prime}(x)=t_{\mathscr{F}}(x)\cup\{bx=x\}$ to obtain $c\in\mathscr{F}_{\varphi}$ s.t. $c\ll b\ll a$. Then by real rank zero and [FA19, 2.7.5], there exists a projection $p$ s.t. $c\ll p\ll a$, so $p\in\mathscr{P}_{\varphi}$. ∎
@@
Enumerate a sequence $(a_{n})$ in $\mathscr{F}_{\varphi}$. As noted in [FA19, 5.3.10], for all $n,m$, there exists $p_{nm}\in\mathscr{P}_{\varphi}$ s.t. $p_{nm}\leq a_{n}+\frac{1}{m}$. Then by the P-filter property, fix $p\in\mathscr{P}_{\varphi}$, a lower bound for every $p_{nm}$. Then for all $n,m$,
@@
Now define an approximate unit $(e_{\gamma})$ for $L_{\varphi}\cap L_{\varphi}^{*}$. Notice that for each $a\in\mathscr{D}$, $(1-e_{\gamma})a-(1-e_{\gamma})=(e_{\gamma}-1)(1-a)=e_{\gamma}(1-a)-(1-a).$ Now fix $F\Subset Y$, $\epsilon>0$, and for each $a\in F$, choose $a^{\prime}\in\mathscr{D}$ s.t. $\|a^{\prime}-a\|<\frac{\epsilon}{2}$. There exists $\gamma$ s.t. for all $a\in F$, $\|e_{\gamma}(1-a^{\prime})-(1-a^{\prime})\|<\frac{\epsilon}{2}$. Therefore
@@
so $(1-e_{\gamma})$ satisfies $t(x)$, so it is realized by $b\in\mathscr{F}_{\varphi}$ s.t. $b\ll a$ for all $a\in Y$. Since $Y$ was arbitrary, we are done.
For $(1)\Rightarrow(3)$, fix countable $Y\subseteq\mathscr{P}_{\varphi}\subseteq\mathscr{F}_{\varphi}$. Then by (2), fix a lower bound $b\in\mathscr{F}_{\varphi}$ for $Y$. Then by lemma 2.6 there exists $p\in\mathscr{P}_{\varphi}$ s.t. $p\ll b$, so $p$ a lower bound for $Y$.
@@
For $(5)\Rightarrow(4)$, enumerate $(p_{n})$ in $\mathscr{P}_{\varphi}$. Then define $a\in C_{+,1}$ by
@@
so $a\in\mathscr{F}_{\varphi}$. By cofinality, fix some $p\in\mathscr{P}_{\varphi}$ s.t. $p\ll a$. Then
@@
Let $\varphi$ be a pure state s.t. $(C,\varphi)$ countably degree-$1$ saturated. Then $\varphi$ has P-excision. If in addition $C$ has real rank zero, then the excising element can always be chosen to be a projection.
@@
By the linearity of $\varphi$, for $\rho_{j}=\varphi$, we can replace $P_{j}(\bar{x})$ by $P_{j}(\bar{x})-d_{j}$, where $d_{j}\in A$ s.t. $\varphi(d_{j})=s_{j}$. Thus, without loss of generality, when $\rho_{j}=\varphi$, assume $s_{j}=0$. Now fix, for each $m\in\mathbb{N}$, $\bar{b}_{m}$ s.t. $\max_{j\leq m}|\rho_{j}(P_{j}(\bar{b}_{m}))-s_{j}|<\frac{1}{m}.$ Now let $X$ be the collection of entries of each $\bar{b}_{m}$ and coefficients of each $P_{j}$. Then $X$ is countable, so $B=C^{*}(X)$ is separable. Mimicking the proof of 2.9, there exists a single $a\in\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. Now define the modified *-polynomials
@@ 3. Infinite Transitivity
Since $t(\bar{x})$ is satisfiable, fix $\bar{b}_{m}$ s.t. $\max_{j\leq m}d(\rho_{j}(P_{j}(\bar{b}_{m})),K_{j})\leq\frac{1}{m}$. Then fix an ultrafilter $\mathscr{U}$ on $\mathbb{N}$. Let $s_{j}=\lim_{m\rightarrow\mathscr{U}}\rho_{j}(P_{j}(\bar{b}_{m}))$. Then
@@
is satisfiable type whose realizations realize $t(\bar{x})$. Now for each $\varphi$, choose $a_{\varphi}\in\mathscr{F}_{\varphi}$ excising the $C^{*}$-subalgebra $B$ generated by coefficients of $P_{j}(\bar{x})$ and entries of $\bar{b}_{m}$. Then, for each $j$ s.t. $\rho_{j}=\varphi$, choose $d_{j}\in B$ s.t. $\varphi(d_{j})=s_{j}$. Define
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and enumerate an orthonormal basis $(\xi_{i})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\bigoplus_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{i})$. For each $i,j\in\mathbb{N}$, lift $a_{ij}+L_{\varphi_{j}}=(\xi_{i})_{j}$ and $b_{ij}+L_{\varphi_{j}}=(T\xi_{i})_{j}$. Define the type
$t(x)=\{\|x\|=\|T\|,\|xa_{ij}-b_{ij}\|^{\varphi_{j}}_{2}=0:i,j\in\mathbb{N}\}.$
Since $t(x)$ consists of conditions from 3.1, it suffices to show that it is satisfiable. Fix $m\in\mathbb{N}$. Because $\{\xi_{j}:j\leq m\}$ is orthogonal, by the Glimm-Kadison theorem, for each $\epsilon>0$ there exists $d\in C$ s.t. $\pi_{j}(d)(\xi_{i})_{j}=(T\xi_{i})_{j}$ for all $i,j\leq m$ and $\|d\|<\|T\|+\epsilon$. That is, $\|da_{ij}-b_{ij}\|^{\varphi}_{2}=0,$ for all $j\leq m$ and $\|d\|<\|T\|+\epsilon$. Therefore $t(x)$ is satisfiable and is realized by $c\in C$, the promised element.
@@
For unitary, notice that by the Borel functional calculus, $T$ has self-adjoint complex logarithm $S\in\bigoplus_{j}\mathscr{B}(K_{j})$. Then, by the self-adjoint case, there exists self-adjoint $s\in C$ so that $P_{K}\pi(s)P_{K}=S$, so $P_{K}\pi(e^{is})P_{K}=T$, so $c=e^{is}$ is our desired unitary. ∎
@@ 4. Existence of P-states
The conclusion of theorem 2.8 can be leveraged to show that, in the right model of ZFC, there are no P-states on many corona algebras. However, in any model of ZFC which admits a P-point (an extremely weak assumption), we can produce a P-state on these coronas.
@@
Given nonunital $A$ with a $\sigma$-unit of projections $(e_{n})$, define the map $p:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p_{X}$ equal to the strict limit of the sequence
$p_{X,m}=\sum_{n\in X,n\leq m}f_{n}$
where $f_{n}=e_{n}-e_{n-1}$, $e_{-1}=0$. For each $\varphi\in S(\mathscr{Q}(A))$, define
@@
The following lemma was originally proven for the Calkin algebra by Zamora-Aviles [ZA10, 2.5.10]. Here we provide a version generalized to coronas with $\sigma$-units of projections.
@@
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra and $(e_{n})$ a $\sigma$-unit of projections. For any decreasing sequence $X_{n}\subseteq\mathbb{N}$ and $a\in\mathscr{Q}(A)_{+,1}$, $a\leq\pi(p_{X_{n}})$ for all $n$ if and only if there exists a pseudointersection $X\subseteq\mathbb{N}$ of $(X_{n})$ s.t. $a\leq\pi(p_{X})$.
@@
Without loss of generality, suppose $X_{1}=\mathbb{N}$ and $X_{n}\setminus X_{n+1}$ infinite. Enumerate $X_{n}\setminus X_{n+1}=\{j^{n}_{i}:i\in\mathbb{N}\}$. Now define $p^{n}:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p^{n}_{Y}=p_{\{j^{n}_{i}:i\in Y\}}$. Since $a=\pi(\tilde{a})\leq\pi(p_{X_{n}})$, we have
$\pi(p_{X_{n}\setminus X_{n+1}}\tilde{a})=(\pi(p_{X_{n}})-\pi(p_{X_{n+1}}))a=a-a=0,$
so $\lim_{k\rightarrow\infty}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|=\lim_{k\rightarrow\infty}\|p^{n}_{\mathbb{N}\setminus k}\tilde{a}\|=0,$ so the function $f_{\epsilon}\in\mathbb{N}^{\mathbb{N}}$ by
@@
is well-defined. Fix $f\in\mathbb{N}^{\mathbb{N}}$ s.t. $f_{\frac{1}{m}}\leq^{*}f$ for all $m$. That is, $f_{\frac{1}{m}}(n)\leq f(n)$ for e.a. $n$. Now we claim that $X=\bigcup_{n\in\mathbb{N}}\{j^{n}_{i}:i\leq f(n)\}$ is our desired pseudointersection. Since $X_{n}$ decreases from $\mathbb{N}$, $X_{n}\setminus X_{n+1}$ partitions $\mathbb{N}$, so
@@
$|X\setminus X_{n}|=\sum_{j=1}^{n-1}(f(j)+1)<\infty,$
thus $X$ is a pseudointersection. Now we claim $a\leq\pi(p_{X})$, and it suffices to show that $\|(1-e_{k})p_{\mathbb{N}\setminus X}\tilde{a}\|\rightarrow 0$. By $f_{\frac{1}{m}}\leq^{*}f$, we have
@@
$\leq\sum_{n=1}^{N-1}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|+\frac{1}{m2^{N}}$
$\leq\frac{1}{m}+\frac{1}{m}\leq\frac{2}{m}$
@@
Let $A$ be a nonunital $\sigma$-unital $C^{*}$-algebra. For every $\sigma$-unit of projections $(e_{n})$ of $A$, there exists $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ such that $\mathscr{X}_{\varphi}=\mathscr{U}$ and $\varphi_{\mathscr{U}}$ is a P-state if and only if $\mathscr{U}$ is a P-point.

### v4

@@ 1.2. Outline
Let $A$ be a nonunital, $\sigma$-unital $C^{*}$-algebra, and $\pi:\mathscr{Q}(A)\rightarrow\mathscr{B}(H)$ a direct sum of countably many inequivalent P-representations $\pi_{j}:\mathscr{Q}(A)\rightarrow\mathscr{B}(H_{j})$ and $K_{j}$ a separable subspace of $H_{j}$. Then the map $\mathscr{Q}(A)\rightarrow\prod_{j}\mathscr{B}(K_{j})$ by $c\mapsto P_{K}\pi(c)P_{K}$ is onto, where $K=\bigoplus_{j}K_{j}$.
As in theorem 3, we can choose $c$ as we like. With the theory complete, under the existence of a P-point, we produce a P-state, on the corona algebra $\mathscr{Q}(A)$ of various nonunital, $\sigma$-unital $A$. However, as shown in [BI13], there is a forcing in which these states cannot exist.
@@ 2.2. Excision, P-Excision, and Quantum P-points
A P-filter is a preordered set $(\mathscr{F},<)$ which is upwards closed and $\sigma$-downwards directed. A P-point is an ultrafilter which is a P-filter under $\subseteq^{*}$, inclusion modulo finite sets. A quantum P-point is a maximal quantum filter which is P-filter under $\ll$. That is, for any countable family $\mathscr{F}\subseteq\mathscr{F}_{\varphi}$, there is $a\in\mathscr{F}_{\varphi}$ such that $a\ll b$ for all $b\in\mathscr{F}$. We call $\varphi$ a P-state if $\mathscr{F}_{\varphi}$ is a quantum P-point. Likewise, we call a representation $\pi$ of $C$ a P-representation if it is equivalent to the GNS-representation corresponding to a P-state.
@@ 2.3. Equivalences
A preordered set $(\mathscr{F},<)$ is countably closed if every decreasing sequence in $\mathscr{F}$ has a lower bound. A subset $\mathscr{P}$ of $\mathscr{F}$ is cofinal if for every $a\in\mathscr{F}$, there exists $b\in\mathscr{P}$ such that $b\leq a$. Lastly, a totally ordered set $(\mathscr{P},<)$ with no endpoints is an $\eta_{1}$-set if for all countable $X,Y\subseteq\mathscr{P}$ such that $a<b$ for all $a\in X$, $b\in Y$, there exists $c\in\mathscr{P}$ such that $a<c<b$ for all $a\in X$, $b\in Y$.
@@
If $C$ has real rank zero, $\varphi$ is a state on $C$, and $(C,\varphi)$ is countably degree-$1$ saturated, then $\mathscr{P}_{\varphi}$ is cofinal in $\mathscr{F}_{\varphi}$.
@@
By [FA19, 5.3.10], for each $\epsilon\in(0,1)$, there exists $q\in\mathscr{P}_{\varphi}$ such that $q\leq a+\epsilon$, so $1-a\leq 1-q+\epsilon$ and
$q(1-a)q\leq q(1-q+\epsilon)q=\epsilon q\leq\epsilon.$
Hence
$\|(1-a)^{\frac{1}{2}}q\|^{2}=\|q(1-a)^{\frac{1}{2}}(1-a)^{\frac{1}{2}}q\|=\|q(1-a)q\|\leq\epsilon,$
and
$\|(1-a)q\|=\|(1-a)^{\frac{1}{2}}(1-a)^{\frac{1}{2}}q\|\leq\|(1-a)^{\frac{1}{2}}\|\epsilon^{\frac{1}{2}}\leq\epsilon^{\frac{1}{2}}.$
Therefore $t(x)$ is satisfiable, so realized by $b\in\mathscr{F}_{\varphi}$ such that $b\ll a$. Then we repeat this with $t^{\prime}(x)=t_{\mathscr{F}}(x)\cup\{bx=x\}$ to obtain $c\in\mathscr{F}_{\varphi}$ such that $c\ll b\ll a$. Then by real rank zero and [FA19, 2.7.5], there exists a projection $p$ such that $c\ll p\ll a$, so $p\in\mathscr{P}_{\varphi}$. ∎
@@
Enumerate a sequence $(a_{n})$ in $\mathscr{F}_{\varphi}$. As noted in [FA19, 5.3.10], for all $n,m$, there exists $p_{nm}\in\mathscr{P}_{\varphi}$ such that $p_{nm}\leq a_{n}+\frac{1}{m}$. Then by the P-filter property, fix $p\in\mathscr{P}_{\varphi}$, a lower bound for every $p_{nm}$. Then for all $n,m$,
@@
Now define an approximate unit $(e_{\gamma})$ for $L_{\varphi}\cap L_{\varphi}^{*}$. Notice that for each $a\in\mathscr{D}$, $(1-e_{\gamma})a-(1-e_{\gamma})=(e_{\gamma}-1)(1-a)=e_{\gamma}(1-a)-(1-a).$ Now fix $F\Subset Y$, $\epsilon>0$, and for each $a\in F$, choose $a^{\prime}\in\mathscr{D}$ such that $\|a^{\prime}-a\|<\frac{\epsilon}{2}$. There exists $\gamma$ such that for all $a\in F$, $\|e_{\gamma}(1-a^{\prime})-(1-a^{\prime})\|<\frac{\epsilon}{2}$. Therefore
@@
so $(1-e_{\gamma})$ satisfies $t(x)$, so it is realized by $b\in\mathscr{F}_{\varphi}$ such that $b\ll a$ for all $a\in Y$. Since $Y$ was arbitrary, we are done.
For $(1)\Rightarrow(3)$, fix countable $Y\subseteq\mathscr{P}_{\varphi}\subseteq\mathscr{F}_{\varphi}$. Then by (2), fix a lower bound $b\in\mathscr{F}_{\varphi}$ for $Y$. Then by lemma 2.6 there exists $p\in\mathscr{P}_{\varphi}$ such that $p\ll b$, so $p$ a lower bound for $Y$.
@@
For $(5)\Rightarrow(4)$, enumerate decreasing $(p_{n})$ in $\mathscr{P}_{\varphi}$. Then define $a\in C_{+,1}$ by
@@
so $a\in\mathscr{F}_{\varphi}$. By cofinality, fix some $p\in\mathscr{P}_{\varphi}$ such that $p\ll a$. Then
@@
Let $\varphi$ be a pure state such that $(C,\varphi)$ countably degree-$1$ saturated. Then $\varphi$ has P-excision. If in addition $C$ has real rank zero, then the excising element can always be chosen to be a projection.
@@
By the linearity of $\varphi$, for $\rho_{j}=\varphi$, we can replace $P_{j}(\bar{x})$ by $P_{j}(\bar{x})-d_{j}$, where $d_{j}\in A$ such that $\varphi(d_{j})=s_{j}$. Thus, without loss of generality, when $\rho_{j}=\varphi$, assume $s_{j}=0$. Now fix, for each $m\in\mathbb{N}$, $\bar{b}_{m}$ such that $\max_{j\leq m}|\rho_{j}(P_{j}(\bar{b}_{m}))-s_{j}|<\frac{1}{m}.$ Now let $X$ be the collection of entries of each $\bar{b}_{m}$ and coefficients of each $P_{j}$. Then $X$ is countable, so $B=C^{*}(X)$ is separable. Mimicking the proof of 2.9, there exists a single $a\in\mathscr{F}_{\varphi}$ which excises $\varphi$ on $B$. Now define the modified *-polynomials
@@ 3. Infinite Transitivity
Since $t(\bar{x})$ is satisfiable, fix $\bar{b}_{m}$ such that $\max_{j\leq m}d(\rho_{j}(P_{j}(\bar{b}_{m})),K_{j})\leq\frac{1}{m}$. Then fix an ultrafilter $\mathscr{U}$ on $\mathbb{N}$. Let $s_{j}=\lim_{m\rightarrow\mathscr{U}}\rho_{j}(P_{j}(\bar{b}_{m}))$. Then
@@
is satisfiable type whose realizations realize $t(\bar{x})$. Now for each $\varphi$, choose $a_{\varphi}\in\mathscr{F}_{\varphi}$ excising the $C^{*}$-subalgebra $B$ generated by coefficients of $P_{j}(\bar{x})$ and entries of $\bar{b}_{m}$. Then, for each $j$ such that $\rho_{j}=\varphi$, choose $d_{j}\in B$ such that $\varphi(d_{j})=s_{j}$. Define
@@
Let $K_{j}$ be a separable subspace of $H_{j}$ and enumerate an orthonormal basis $(\xi_{i})$ of $K=\bigoplus_{j}K_{j}$ compatible with each $K_{j}$. Every $T\in\prod_{j}\mathscr{B}(K_{j})$ is determined by $(T\xi_{i})$. For each $i,j\in\mathbb{N}$, lift $a_{ij}+L_{\varphi_{j}}=(\xi_{i})_{j}$ and $b_{ij}+L_{\varphi_{j}}=(T\xi_{i})_{j}$. Define the type
$t(x)=\{\|x\|\leq\|T\|,\|xa_{ij}-b_{ij}\|^{\varphi_{j}}_{2}=0:i,j\in\mathbb{N}\}.$
Since $t(x)$ consists of conditions from 3.1, it suffices to show that it is satisfiable. Fix $m\in\mathbb{N}$. Because $\{\xi_{j}:j\leq m\}$ is orthogonal, by the Glimm-Kadison theorem, for each $\epsilon>0$ there exists $d\in C$ such that $\pi_{j}(d)\xi_{i}=T\xi_{i}$ for all $i,j\leq m$ and $\|d\|<\|T\|+\epsilon$. That is, $\|da_{ij}-b_{ij}\|^{\varphi_{j}}_{2}=0,$ for all $j\leq m$ and $\|d\|<\|T\|+\epsilon$. Therefore $t(x)$ is satisfiable and is realized by $c\in C$. Finally, $\|c\|\geq\|\pi(c)\|\geq\|T\|$, so $\|c\|=\|T\|$ and $P_{K}\pi(c)P_{k}=T$.
@@
For unitary, notice that by the Borel functional calculus, $T$ has self-adjoint complex logarithm $S\in\prod_{j}\mathscr{B}(K_{j})$. Then, by the self-adjoint case, there exists self-adjoint $s\in C$ so that $P_{K}\pi(s)P_{K}=S$, so $P_{K}\pi(e^{is})P_{K}=T$, so $c=e^{is}$ is our desired unitary. ∎
@@ 4. Existence of P-states
The conclusion of theorem 2.8 can be leveraged to show it is relatively consistent with ZFC that there are no P-states on many corona algebras. However, in any model of ZFC which admits a P-point (an extremely weak assumption), we can produce a P-state on many coronas. To accomplish this, we want to mimic the construction of diagonal pure states on $\mathscr{Q}(H)$. As such, we need a generalization of an orthonormal sequence of vectors.
@@
Let $A$ be nonunital, $\sigma$-unital. We call a sequence $(q_{j})$ of projections a sub-$\sigma$ sequence of $A$ if there is a $\sigma$-unit $(e_{j})$ of $A$ such that $q_{j}\leq e_{j}-e_{j-1}$, $e_{-1}=0$. For a sub-$\sigma$ sequence, define the map $p:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p_{X}$ equal to the strict limit of the sequence
$p_{X,n}=\sum_{\begin{subarray}{c}j\in X\\ j\leq n\end{subarray}}q_{j}$
and we abuse the notation $p_{X}=\sum_{j\in X}q_{j}$. For each $\varphi\in S(\mathscr{Q}(A))$, define
@@
We call a sub-$\sigma$ sequence scalar if $q_{j}$ is a scalar projection for all $j$ and there is a faithful, nondegenerate representation $\rho:A\rightarrow\mathscr{B}(H)$ such that $\rho(q_{j})$ has finite rank for all $j$.
To justify the definition of a scalar sub-$\sigma$ sequence, let $A=\bigoplus_{i\in I}\mathscr{K}(H_{i})$ such that $\bigoplus_{i\in I}H_{i}$ is infinite dimensional, separable. Then $A$ has a scalar sub-$\sigma$ sequence. Namely, we identify $A\subseteq\mathscr{B}(\bigoplus_{i}H_{i})$, fix an orthonormal basis $(\xi_{j})$ of $\bigoplus_{i}H_{i}$ compatible with the direct sum, making $q_{j}=\text{proj}_{\mathbb{C}\xi_{j}}$ scalar sub-$\sigma$. Any subsequence of a scalar sub-$\sigma$ sequence is still scalar sub-$\sigma$.
The following lemma was originally proven for the Calkin algebra by Zamora-Avilés [ZA10, 2.5.10]. Here we provide a version generalized to coronas of algebras with a sub-$\sigma$ sequence.
@@
Let $A$ be a nonunital, $\sigma$-unital $C^{*}$-algebra. For every sub-$\sigma$ sequence in $A$, decreasing sequence $X_{n}\subseteq\mathbb{N}$, and $a\in\mathscr{Q}(A)_{+,1}$, we have $a\ll\pi(p_{X_{n}})$ for all $n$ if and only if there exists a pseudointersection $X\subseteq\mathbb{N}$ of $(X_{n})$ such that $a\ll\pi(p_{X})$.
@@
Without loss of generality, suppose $X_{1}=\mathbb{N}$ and $X_{n}\setminus X_{n+1}$ infinite. Enumerate $X_{n}\setminus X_{n+1}=\{j^{n}_{i}:i\in\mathbb{N}\}$. Now define $p^{n}:\mathscr{P}(\mathbb{N})\rightarrow\mathrm{Proj}(\mathscr{M}(A))$ by $p^{n}_{Y}=p_{\{j^{n}_{i}:i\in Y\}}$. Fix a lift $\tilde{a}\in\mathscr{M}(A)$, so $a=\pi(\tilde{a})\ll\pi(p_{X_{n}})$, and we have
$\pi(p_{X_{n}\setminus X_{n+1}}\tilde{a})=(\pi(p_{X_{n}})-\pi(p_{X_{n+1}}))a=a-a=0.$
By the sub-$\sigma$ property, for each $m$, there is $k$ such that $(1-e_{k})p^{n}_{\mathbb{N}}=p^{n}_{\mathbb{N}\setminus m}$. This yields
$\lim_{k\rightarrow\infty}\|p^{n}_{\mathbb{N}\setminus k}\tilde{a}\|=\lim_{k\rightarrow\infty}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|=0,$
so the function $f_{\epsilon}\in\mathbb{N}^{\mathbb{N}}$ by
@@
is well-defined. Fix $f\in\mathbb{N}^{\mathbb{N}}$ such that $f_{\frac{1}{m}}\leq^{*}f$ for all $m$. That is, $f_{\frac{1}{m}}(n)\leq f(n)$ for eventually all $n$. Now we claim that $X=\bigcup_{n\in\mathbb{N}}\{j^{n}_{i}:i\leq f(n)\}$ is our desired pseudointersection. Since $X_{n}$ decreases from $\mathbb{N}$, $X_{n}\setminus X_{n+1}$ partitions $\mathbb{N}$, so
@@
$|X\setminus X_{n}|=\sum_{j=1}^{n-1}f(j)<\infty,$
thus $X$ is a pseudointersection. Now we claim $a\ll\pi(p_{X})$, and it suffices to show that $\|(1-e_{k})p_{\mathbb{N}\setminus X}\tilde{a}\|\rightarrow 0$. By $f_{\frac{1}{m}}\leq^{*}f$, we have
@@
$\leq\sum_{n=1}^{N-1}\|(1-e_{k})p_{X_{n}\setminus X_{n+1}}\tilde{a}\|+\frac{1}{m}$
$\leq\frac{1}{m}+\frac{1}{m}=\frac{2}{m}$
@@
This lemma shows us that P-states cannot naturally arise from non-P-points.
Corollary 4.3.
Let $A$ be nonunital, $\sigma$-unital. For every sub-$\sigma$ sequence, there exists a natural $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ such that $\mathscr{X}_{\varphi}=\mathscr{U}$ and $\varphi_{\mathscr{U}}$ is a P-state only if $\mathscr{U}$ is a P-point.
Proof.
[…]
Suppose $\mathscr{U}$ is not a P-point, and fix decreasing $X_{n}\in\mathscr{U}$ with no pseudointersection in $\mathscr{U}$. Our claim is that $\pi(p_{X_{n}})$ has no lower bound in $\mathscr{F}_{\varphi_{\mathscr{U}}}$. By 4.2, any lower bound $a\in\mathscr{F}_{\varphi}$ of $(\pi(p_{X_{n}}))$ must in fact have $a\ll\pi(p_{X})$, where $X$ is a pseudointersection of $(X_{n})$, therefore no such $a$ exists. ∎
@@
Let $A$ be nonunital, $\sigma$-unital. For every scalar sub-$\sigma$ sequence in $A$, there exists a canonical $\varphi:\beta\mathbb{N}\setminus\mathbb{N}\hookrightarrow P(\mathscr{Q}(A))$ such that $\mathscr{X}_{\varphi}=\mathscr{U}$ and $\varphi_{\mathscr{U}}$ is a P-state if and only if $\mathscr{U}$ is a P-point. In particular, if $\mathscr{U}$ is a P-point, lower bounds of countable families in $\mathscr{F}_{\varphi_{\mathscr{U}}}$ can be chosen of the form $\pi(p_{X})$ for $X\in\mathscr{U}$.